# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v3

@@ II Related Works
Chao et al. [chao2024systematic] 2024
Vulnerabilities, Attacks, Defenses, and Formalization
in Anonymity Networks.
@@ III-A Tor Architecture
Encryption Layer Application layer (TLS)
Network/transport layer (IPsec,
OpenVPN, WireGuard)
Transport layer (over UDP)
Application layer (TLS) +
multiple network layers
Encryption Scope
Only app data (e.g., web
content)
Entire network traffic App data with transport layer
Multi-layered encryption of
full paths
IP Address Hidden? No
Yes from websites
No from VPN provider
No Yes
Anonymity No
Partial (VPN sees
identity and activity)
No Yes
@@
DNS Leakage Protection No
Yes, if DNS goes
through VPN
No
Yes, DNS requests
resolved within Tor
Resistant to Censorship?
Partially ( depends on
SNI and DoH
Depends on obfuscation
layer used
No Strong resistance
Packet Size
$\sim$1,300 bytes
(TCP + TLS payload)
1,400–1,500 bytes
(depends on protocol)
1,200–1,400 bytes
(UDP + QUIC)
Fixed-size
(512- or 514-byte cells)
@@
Tor applies a layered encapsulation and segmentation process to ensure secure and reliable network packet delivery. Specifically, Figure 3 illustrates the data representation across different layers of transport within the Tor network. At the application layer, Tor encapsulates data into fixed-size 512- or 514-byte cells, which serve as the fundamental unit of communication. These cells are subsequently encrypted and aggregated into TLS records, with each record typically containing several complete Tor cells. As TLS records traverse the network stack, they are further segmented into multiple TCP packets, constrained by the maximum segment size (MSS) of 1448 bytes for Ethernet. In the Tor network, each circuit can carry multiple streams, with each stream corresponding to a separate TCP connection. To manage flow control, Tor employs SENDME cells: one is sent for every 50 incoming cells per stream and one for every 100 incoming cells per circuit. In this way, Tor can produce traffic patterns that reflect the substantial data volumes generated by modern websites [panchenko2016website, deng2025countmamba].
@@ III-B Tor Traffic Features
Cell size Websites differ in total packet/cell size due to varying content. Packet/cell size. Excluded from WF defenses because of Tor’s fixed-size cells.
@@ III-B2 Traffic Length
III-B3 Cell Size
In Tor, traffic is encapsulated into fixed-size units called cells (typically 512 or 514 bytes), so attackers cannot directly infer information from individual cells [wang2016realistically]. In early studies, researchers captured Tor traffic in a closed-world setting, where the total cell count and packet length for each webpage were known. Thus, cell size is widely used in traditional ML-based WF attacks [wang2014effective, cai2012touching, wang2013improved, zhuo2017website, cui2019revisiting]. However, in real-world scenarios, attackers cannot precisely measure the total traffic length because the start and end Tor cells are unknown. Consequently, recent website fingerprinting studies seldom rely on cell size for website classification.
@@ III-D2 Evaluation metrics
Bandwidth overhead / Data overhead [gong2020zero, lu2023lightweight]: The data overhead $O(D)$ of defense method $D$ on $P$ is defined as the ratio of the total amount of dummy data to the total amount of real data. $P$ denotes the original trace. $P^{{}^{\prime}}$ denotes the trace after implementing some defense $D$.
$O(D)=\frac{\left|P^{{}^{\prime}}\right|-\left|P\right|}{\left|P\right|}$ (10)
@@ IV-A2 Open-world dataset
Dataset Name Year Dataset Size Features
Included
Multi-tab
@@
Traffic timing, Packet length, Traffic direction,
Packet ordering, Traffic bursts
CUMUL [panchenko2016website] 2016 211,148 118,884
Traffic timing, Cell size, Packet length,
Traffic direction, Traffic ordering, Traffic bursts
k-fingerprinting [hayes2016k] 2016
@@
NA
Traffic timing, Packet length, Traffic direction,
Traffic ordering, Traffic bursts
[…]
Traffic timing, Packet length, Traffic direction,
Traffic ordering, Traffic bursts
[…]
Traffic timing, Cell size, Packet length,
Traffic direction, Traffic ordering, Traffic bursts
DS-19 [gong2020zero] 2020 $\times100$ $00×\times1$
Traffic timing, Cell size, Packet length,
Traffic direction, Traffic ordering, Traffic bursts
Tik-Tok/W-T [rahman2019tik] 2019 $0×\times1$ $00×\times1$
Traffic timing, Packet length, Traffic direction,
Traffic ordering, Traffic bursts
ARES [deng2023robust] 2023 $\times1,00$0 $\times1,00$0
Traffic timing, Traffic direction, Traffic ordering,
Traffic bursts
[…]
Traffic timing, Cell size, Packet length,
Traffic direction, Traffic ordering, Traffic bursts
TMWF [jin2023transformer] 2023
[…]
$\text{Chrome}_{M}$: $00×\times1$
[…]
$\text{TBB}_{M}$: $00×\times1$
Traffic timing, Traffic direction, Traffic ordering,
Traffic bursts
@@ V-A Survey on WF attacks
Machine Learning
Naïve Bayes  [herrmann2009website]
@@
BalanceCascade-
XGBoost
 [xu2018multi]
@@
Deep Learning
CNN  [hong2024website, shen2023subverting, bhat2018var, yuan2025mw3f, gong2024wfcat, cui2020more]
@@
Few-shot Learning  [zou2022efficient, hong2024website, meng2025beyond]
@@ V-A1 Traditional ML-based WF Attack Models
Research Year Architecture
Dataset
Closed World Open World
@@
Wang et al. [wang2013improved] 2013
SVM and OSAD
CS-BUFLO [cai2012touching]
@@
90.00 NA NA
[…]
Wang et al. [wang2016realistically] 2016
Splitting algorithm
and K-NN
[…]
Splitting algorithm
and Time-KNN
[…]
Splitting algorithm
and SVM
@@ V-A2 DL-based WF Attack models
TinyWFP [shen2023subverting]
Wavelet transform and
Lightweight Neural Network
AWF [rimmer2017automated]
DF [sirinam2018deep]
2,500
1,000
99.20
98.10
20,000 97.50 4.50
WFCAT [gong2024wfcat]
IAT histogram
and CNN
WFCAT [gong2024wfcat] $\times100$ 94.47 NA NA NA
CF [xie2024contrastive]
Contrastive learning
and Few-shot
Wang14 [wang2014effective]
AWF [rimmer2017automated]
DF [sirinam2018deep]
@@
SMC [tan2024adaptability]
Mixed ( [wang2014effective]
 [rimmer2017automated, gong2020zero])
59.11 76.20 80.70 82.37 82.61 78.96 83.69 NA NA
@@
Concept drift in website fingerprinting can occur when websites modify their structure or content. Even small changes, such as adding images, updating ads, or reorganizing page elements, alter the resulting traffic patterns. These shifts cause the distribution of features like packet sizes, timing, and traffic length to deviate from the data on which models were trained. To address the problem of concept drift, Ding et al. [ding2024multi] presented LRCT, a DL-based WF attack model designed to improve classification accuracy and efficiency while addressing concept drift—a phenomenon where model performance degrades over time due to network changes. The LRCT network effectively leverages the temporal learning advantages of Local RNN and the spatial learning strengths of CNN by designing the local feature extraction block (denoted as LRC Block). The network then uses a Transformer Encoder to capture more robust global features from the low-dimensional data. Experimental results comparing LRCT with state-of-the-art WF methods demonstrate that the proposed LRCT achieves 99.34% accuracy in a closed-world scenario, outperforming other benchmark models.
@@ V-A3 Multi-tab WF attacks
Research Architecture
Dataset
Name
Closed World
Precision(%)
Open World
Precision(%)
@@
BAPM [guan2021bapm]
Self-attention
and CNN
Xu et al. [xu2018multi]
and Manually
Merged
66.80 77.30 NA NA 84.90 NA NA NA
TMWF [jin2023transformer]
Transformer
and CNN
BAPM [guan2021bapm],
Walkie-Talkie [wang2017walkie]
and Manually
Merged
95.50 NA NA NA >>70.00 $\sim$70.00 >>70.00 >>70.00
ARES [deng2023robust]
Top-k multi-head
self-attention and CNN
ARES [deng2023robust] 91.30 91.30 90.20 81.80 75.20 58.70 52.80 43.80
FingerMamba [liu2024fingermamba]
Mamba model
and CNN
ARES [deng2023robust]
and Manually
Merged
83.00 80.80 79.70 75.70 81.00 79.00 >>75.00 >>75.00
MTWFAs [yin2024traces] FCN
DS-19 [gong2020zero],
ARES [deng2023robust]
and Manually
Merged
68.60 68.60 70.00 72.10 73.30 74.80 73.90 71.80
FMWF [meng2025beyond] Few-shotFine-tuning
ARES [deng2023robust]
and Manually
Merged
88.50 77.90 89.00 84.90 NA NA NA NA
MW3F [yuan2025mw3f]
Transformer
and CNN
ARES [deng2023robust],
Deng et al. [deng2024robust],
and Manually
Merged
92.91 93.11 92.66 89.88 >>90.00 >>90.00 >>90.00 >>90.00
CountMamba [deng2025countmamba]
State Space Model
and CNN
ARES [deng2023robust] and
TMWF [jin2023transformer]
87.33 81.52 81.26 73.89 85.02 81.17 79.98 75.60
@@ V-B1 Adaptive Padding-Based Defenses
WF Attack Model
Closed World
Acc (%)
Open World
F1 (%)
[…]
1) No changes to anonymization protocols required; 2) Obfuscation through parallel page loading.
1) High bandwidth overhead; 2) In closed-world assumption for all web pages.
[…]
1) Adaptive padding based on traffic bursts; 2) Lightweight protection with zero communication latency; 3) Low overhead and deployed in real-world Tor settings.
1) Less effective against deep learning-based and adaptive attacks; 2) Less effective in open-world scenarios; 3) Zero-latency but approximate high bandwidth overhead; 4) Do not significantly alter the burst patterns.
@@
1) High effectiveness with lightweight bandwidth use; 2) Effective obfuscation of trace fronts; 3) Zero latency overhead; 4) Randomizes dummy packet at the beginning of a page load which will reduce the latency overhead.
1) Overhead is dataset dependent; 2) Require timestamp information from original traffic samples.
@@
1) Low time delay; 2) Randomized bidirectional padding.
Higher bandwidth overhead than some lightweight methods.
@@
1) Low-latency; 2) Controllable overhead.
1) Feature leakage from the timing trace remains; 2) Low-efficiency in DL-based WF attack models.
@@
1) Lightweight and zero latency; 2) Beta distribution-based padding and conceals critical features; 3) Higher randomness and flexibility.
1) Limited scope to packet injection; 2) Low-efficiency defense against DL-based WF attack models.
@@ V-B2 Regularization-based defenses
WF Attack Model
Closed World
Acc (%)
[…]
Resistant to fine-grained classifiers.
1) Extremely high bandwidth overhead and high latency; 2) Leaks coarse-grained features.
[…]
1) Congestion sensitivity; 2)Rate adaptation with limited leakage; 3) Adjustable overhead.
1) High bandwidth overhead; 2) Client-Server coordination required.
[…]
1) No pre-established database of website traffic patterns required; 2) Real-time operation.
1) Impractical for real-world deployment.; 2) High bandwidth and latency overhead.
[…]
1) Obfuscates session boundaries that an attacker is impossible to obtain the beginning and end times of the trace.; 2) Resistant to the ”Split Problem”.
1) User behavior assumptions required; 2) Less effective if there is a long interval between two opening websites.
[…]
1) No packet content or full trace knowledge required; 2) Adjustable overhead; 3) Achieves the lowest attack accuracy among all methods with comparable overhead.
1) Requires parameter tuning and delaying data transmission; 2) Only focuses on regulating the coarse feature.
[…]
1) Provides strong protection against DL-based WF attacks; 2) Effective against adversarial training; 3) Less information leakage.
1) High bandwidth overhead; 2) Extra buffer required for each traffic regulation; 3) May cause a high traffic delay due to the traffic regulation.
[…]
Later, Lu et al. [lu2018dynaflow] introduced DynaFlow, a WF defense mechanism by dynamically adjusting website traffic flows using the average inter-packet arrival time from the previous period. DynaFlow enhances security while improving efficiency by over 40% compared to previous WF defense mechanisms [cai2014systematic, dyer2012peek, cai2014cs], eliminating the need for a pre-established database and extending protection to dynamically-generated websites. Experimental evaluations demonstrate that DynaFlow significantly reduces attacker accuracy while maintaining acceptable latency and bandwidth overheads, making it a promising and practical solution for safeguarding user privacy amid increasing surveillance. To address the challenges of high bandwidth and latency overhead, as well as the reliance on additional infrastructure, Holland et al. [holland2020regulator] proposed RegularTor. RegularTor mitigates these attacks by leveraging common patterns in typical web browsing traffic, thereby reducing both the resource overhead of the defense and the accuracy of state-of-the-art WF classifiers. Experimental results demonstrate that RegulaTor reduces the accuracy of the Tik-Tok attack from 66% to 25.4%, significantly outperforming prior defense methods. To effectively defend against deep learning-based WF attacks with lower overhead, Chen et al. [chen2023web] proposed Timing-Assistant Traffic Burst Shaping (BM). This approach addresses the limitations of existing defenses, such as WTF-PAD and Walkie-Talkie, by incorporating temporal features. BM redefines traffic bursts using dynamic time thresholds and merges sequences from sensitive and non-sensitive websites into a unified “super-sequence” to obscure burst-level patterns. Later, Shen et al. [shen2024real] presented Palette, a practical defense against WF attacks that threaten user privacy in Tor. Unlike traditional defenses, Palette focuses on traffic cluster anonymization by grouping websites with similar traffic patterns into clusters and regulating their traffic to conform to a uniform pattern. This approach makes it significantly harder for attackers to distinguish between websites within the same cluster, effectively mitigating the risk posed by modern deep learning-based WF attacks.
Defense Model Effectiveness Strengths Drawbacks
WF Attack Model
Closed World
Acc (%)
Open World
Acc (%)
[…]
Optimal classification strategy.
High computational complexity of SCS.
[…]
1) Efficient decoy selection and storage; 2) Low overhead and half-duplex communication simplifies traffic structure.
1) Requires decoy page priori knowledge about each site; 2) Cannot protect dynamic Content Well; 3) Webpage loading mechanism needs to be modified for the browser and deployment is challenging.
@@
1) Bi-directional burst morphing; 2) Lower Bandwidth Overhead and zero-delay transmission.
1) Computational overhead for initialization phase; 2) Sensitivity to target selection; 3) Limited in defending against DL-based WF attacks.
@@
1) Innovative use of fuzzing and neuron feedback; 2) Minimal overhead with high defense performance.
1) Prior knowledge of website required; 2) Still vulnerable to adversarial training.

### v4

@@ II Related Works
Chao et al. [chao2024systematic] 2024 Vulnerabilities, Attacks, Defenses, and Formalization in Anonymity Networks.
@@ III-A Tor Architecture
Encryption Layer Application layer (TLS) Network/transport layer (IPsec, OpenVPN, WireGuard) Transport layer (over UDP) Application layer (TLS) + multiple network layers
Encryption Scope Only app data (e.g., web content) Entire network traffic App data with transport layer Multi-layered encryption of full paths
IP Address Hidden? No Yes from websites No from VPN provider No Yes
Anonymity No Partial (VPN sees identity and activity) No Yes
@@
DNS Leakage Protection No Yes, if DNS goes through VPN No Yes, DNS requests resolved within Tor
Resistant to Censorship? Partially ( depends on SNI and DoH Depends on obfuscation layer used No Strong resistance
Packet Size $\sim$1,300 bytes (TCP + TLS payload) 1,400–1,500 bytes (depends on protocol) 1,200–1,400 bytes (UDP + QUIC) Fixed-size (512- or 514-byte cells)
@@
Tor applies a layered encapsulation and segmentation process to ensure secure and reliable network packet delivery. Specifically, Figure 3 illustrates the data representation across different layers of transport within the Tor network. At the application layer, Tor encapsulates data into fixed-size 512- or 514-byte cells, which serve as the fundamental unit of communication. These cells are subsequently encrypted and aggregated into TLS records, with each record typically containing several complete Tor cells. As TLS records traverse the network stack, they are further segmented into multiple TCP packets, constrained by the maximum segment size (MSS) of 1460 bytes for Ethernet. In the Tor network, each circuit can carry multiple streams, with each stream corresponding to a separate TCP connection. To manage flow control, Tor employs SENDME cells: one is sent for every 50 incoming cells per stream and one for every 100 incoming cells per circuit. In this way, Tor can produce traffic patterns that reflect the substantial data volumes generated by modern websites [panchenko2016website, deng2025countmamba].
@@ III-B Tor Traffic Features
Cell/Packet size Websites traffic differ in TLS record size, but Tor has a fixed-size Cell. Packet/cell size. Excluded from WF defenses because of Tor’s fixed-size cells.
@@ III-B2 Traffic Length
III-B3 Cell Size / Packet Length
In Tor, traffic is encapsulated into fixed-size units called cells (typically 512 or 514 bytes), so attackers cannot directly infer information from individual cells [wang2016realistically]. In early studies, researchers captured Tor traffic in a closed-world setting, where the total cell count and packet length for each webpage were known. Thus, cell size is widely used in traditional ML-based WF attacks [wang2014effective, cai2012touching, wang2013improved, zhuo2017website, cui2019revisiting]. However, when Tor traffic is captured at the TLS record level, cells appear as variable-size packets rather than fixed-size units. Compared to cell size, packet length information extracted from TLS records can leak more webpage characteristics due to the inherently varying sizes of web content.
@@ III-D2 Evaluation metrics
Bandwidth overhead / Data overhead [gong2020zero, lu2023lightweight]: The bandwidth overhead $B(D)$ of defense method $D$ on $P$ is defined as the ratio of the total amount of dummy data to the total amount of real data. $P$ denotes the original trace. $P^{{}^{\prime}}$ denotes the trace after implementing some defense $D$.
$B(D)=\frac{\left|P^{{}^{\prime}}\right|-\left|P\right|}{\left|P\right|}$ (10)
@@ IV-A2 Open-world dataset
Dataset Name Year Dataset Size Features Included Multi-tab
@@
[…]
CUMUL [panchenko2016website] 2016 211,148 118,884 Traffic timing, Packet length, Traffic direction, Traffic ordering, Traffic bursts
@@
[…]
DS-19 [gong2020zero] 2020 $\times100$ $00×\times1$ Traffic timing, Traffic direction, Traffic ordering, Traffic bursts
Tik-Tok/W-T [rahman2019tik] 2019 $0×\times1$ $00×\times1$ Traffic timing, Traffic direction, Traffic ordering, Traffic bursts
ARES [deng2023robust] 2023 $\times1,00$0 $\times1,00$0 Traffic timing, Traffic direction, Traffic ordering, Traffic bursts
@@ V-A Survey on WF attacks
Machine Learning Naïve Bayes  [herrmann2009website]
@@
BalanceCascade- XGBoost  [xu2018multi]
@@
Deep Learning CNN  [hong2024website, shen2023subverting, bhat2018var, yuan2025mw3f, gong2024wfcat, cui2020more]
@@
Few-shot Learning  [zou2022efficient, hong2024website, meng2025beyond, shen2025swallow]
@@ V-A1 Traditional ML-based WF Attack Models
Research Year Architecture Dataset Closed World Open World
@@ V-A2 DL-based WF Attack models
TinyWFP [shen2023subverting] Wavelet transform and Lightweight Neural Network AWF [rimmer2017automated] DF [sirinam2018deep] 2,500 1,000 99.20 98.10 20,000 97.50 4.50
WFCAT [gong2024wfcat] IAT histogram and CNN WFCAT [gong2024wfcat] $\times100$ 94.47 NA NA NA
@@
SMC [tan2024adaptability] Mixed ( [wang2014effective]  [rimmer2017automated, gong2020zero]) 59.11 76.20 80.70 82.37 82.61 78.96 83.69 NA NA
@@
The SOTA WF attacks either rely on the strong assumption that classifiers are trained and deployed under identical network conditions or suffer significant performance degradation when confronted with WF defenses. To address this limitation, Shen et al. [shen2025swallow] proposed Swallow, a transfer-robust WF attack that rapidly adapts to new network conditions while remaining resilient to a wide range of WF defenses. Specifically, the authors introduced a novel trace representation, termed Consistent Interaction Feature (CIF), which aligns traffic distributions across diverse network conditions to capture invariant features. They further designed three data augmentation strategies to simulate traffic variations induced by different network conditions. Results from both closed-world and open-world evaluations show that Swallow significantly outperforms existing SOTA WF attacks. Notably, with only five labeled instances per website for fine-tuning through few-shot learning, Swallow achieves an average accuracy improvement of 17.50% over prior methods.
@@
Concept drift in website fingerprinting can occur when websites modify their structure or content. Even small changes, such as adding images, updating ads, or reorganizing page elements, alter the resulting traffic patterns. These shifts cause the distribution of features like packet sizes, timing, and traffic length to deviate from the data on which models were trained. To address the problem of concept drift, Ding et al. [ding2024multi] presented LRCT, a DL-based WF attack model designed to improve classification accuracy and efficiency while addressing concept drift, a phenomenon where model performance degrades over time due to network changes. The LRCT network effectively leverages the temporal learning advantages of Local RNN and the spatial learning strengths of CNN by designing the local feature extraction block (denoted as LRC Block). The network then uses a Transformer Encoder to capture more robust global features from the low-dimensional data. Experimental results comparing LRCT with state-of-the-art WF methods demonstrate that the proposed LRCT achieves 99.34% accuracy in a closed-world scenario, outperforming other benchmark models.
@@ V-A3 Multi-tab WF attacks
Research Architecture Dataset Name Closed World Precision(%) Open World Precision(%)
@@
BAPM [guan2021bapm] Self-attention and CNN Xu et al. [xu2018multi] and Manually Merged 66.80 77.30 NA NA 84.90 NA NA NA
TMWF [jin2023transformer] Transformer and CNN BAPM [guan2021bapm], Walkie-Talkie [wang2017walkie] and Manually Merged 95.50 NA NA NA >>70.00 $\sim$70.00 >>70.00 >>70.00
ARES [deng2023robust] Top-k multi-head self-attention and CNN ARES [deng2023robust] 91.30 91.30 90.20 81.80 75.20 58.70 52.80 43.80
FingerMamba [liu2024fingermamba] Mamba model and CNN ARES [deng2023robust] and Manually Merged 83.00 80.80 79.70 75.70 81.00 79.00 >>75.00 >>75.00
MTWFAs [yin2024traces] FCN DS-19 [gong2020zero], ARES [deng2023robust] and Manually Merged 68.60 68.60 70.00 72.10 73.30 74.80 73.90 71.80
FMWF [meng2025beyond] Few-shotFine-tuning ARES [deng2023robust] and Manually Merged 88.50 77.90 89.00 84.90 NA NA NA NA
MW3F [yuan2025mw3f] Transformer and CNN ARES [deng2023robust], Deng et al. [deng2024robust], and Manually Merged 92.91 93.11 92.66 89.88 >>90.00 >>90.00 >>90.00 >>90.00
CountMamba [deng2025countmamba] State Space Model and CNN ARES [deng2023robust] and TMWF [jin2023transformer] 87.33 81.52 81.26 73.89 85.02 81.17 79.98 75.60
@@ V-B1 Adaptive Padding-Based Defenses
WF Attack Model Closed World Acc (%) Open World F1 (%)
@@
[…]
@@
WF Attack Model Closed World Acc (%)
@@
[…]
V-B2 Regularization-based defenses
Later, Lu et al. [lu2018dynaflow] introduced DynaFlow, a WF defense mechanism by dynamically adjusting website traffic flows using the average inter-packet arrival time from the previous period. DynaFlow enhances security while improving efficiency by over 40% compared to previous workflow defense mechanisms [cai2014systematic, dyer2012peek, cai2014cs], eliminating the need for a pre-established database and extending protection to dynamically generated websites. Experimental evaluations demonstrate that DynaFlow significantly reduces attacker accuracy while maintaining acceptable latency and bandwidth overhead, making it a promising and practical solution for safeguarding user privacy in the face of increasing surveillance. To address the challenges of high bandwidth and latency overhead, as well as the reliance on additional infrastructure, Holland et al. [holland2020regulator] proposed RegularTor. RegularTor mitigates these attacks by leveraging common patterns in typical web browsing traffic, thereby reducing both the resource overhead of the defense and the accuracy of state-of-the-art WF classifiers. Experimental results demonstrate that RegulaTor reduces the accuracy of the Tik-Tok attack from 66% to 25.4%, significantly outperforming prior defense methods. To effectively defend against deep learning-based WF attacks with lower overhead, Chen et al. [chen2023web] proposed Timing-Assistant Traffic Burst Shaping (BM). This approach addresses the limitations of existing defenses, such as WTF-PAD and Walkie-Talkie, by incorporating temporal features. BM redefines traffic bursts using dynamic time thresholds and merges sequences from sensitive and non-sensitive websites into a unified “super-sequence” to obscure burst-level patterns. Later, Shen et al. [shen2024real] presented Palette, a practical defense against WF attacks that threaten user privacy in Tor. Unlike traditional defenses, Palette focuses on traffic cluster anonymization by grouping websites with similar traffic patterns into clusters and regulating their traffic to conform to a uniform pattern. This approach makes it significantly harder for attackers to distinguish between websites within the same cluster, effectively mitigating the risk posed by modern deep learning-based WF attacks.
Defense Model Effectiveness Strengths Drawbacks
WF Attack Model Closed World Acc (%) Open World Acc (%)
[…]